OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4 plz help

OpenStudy (anonymous):

f(x) = y y = 3(x-1) x = 4 y = 3(4-1) y = 3(3) y = 9 f(x) = 9 f(4) = 9 . So this defines the point (x,y) = (4,9). The slope-intercept form of the equation is: y = mx+b y = 3(x-1) y = 3x -3 . When x=0 y = -3 . That is the y-intercept

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ur very welcome

OpenStudy (anonymous):

glad i can help